OpenStudy (jotopia34):

REDOX question. I can not get this straight, Which is the [red] and which is the [ox]? SO4(2-) or SO3(2-)? Please explain as well. thx!

OpenStudy (australopithecus):

LEO the lion says GRR Lose electrons oxidation Gain electrons reduction So, the oxidizing agent is reduced the reducing agent is oxidized When negative charge goes up that means that the number of electrons the atom or molecule has, has increased When positive charge goes up that means the number electrons the atom or molecule has, has decreased When the molecule/element is neutral it means that the number of protons equal the number of electrons. When looking at individual atoms for their oxidation state we simply follow two rules: 1. Oxidation state is the charge atoms would have if every bond was a 100% ionic. 2. Which ever is more electronegative gets the electrons so with SO4 each oxygen will have a -2 charge because it is more electronegative than sulphur and with -2 charge it obeys the octet rule (look up periodic table trend electronegativity on google images) so, -2*4 = -8 since the entire compound has a charge of -2, we can set up a simple equation to figure out sulphurs oxiation state -2 = x + (-8) where, x = the oxidation state of sulphur. Therefore Sulphur has an oxidation state of +6 Now look at SO3 same deal this time -2*3 = -6 Set up the equation -2 = x + (-6) therefore the oxidation state of the sulphur is +4 Now just read what I wrote above to see which one was reduced and which was oxidized

OpenStudy (australopithecus):

Sorry for the really long answer I just wanted to be thorough.

OpenStudy (jotopia34):

You may be answering this, but could this be correct? if equn is SO4(2-) + 2H+ +2e = SO3(2-) +H2O is that like saying Ox+ne = Red? and therefore the sulfate ion would be the "ox"?

OpenStudy (australopithecus):

Just read what I wrote and memorize the few things if you want to do well on your test

OpenStudy (jotopia34):

What if Im not given that data?

OpenStudy (jotopia34):

I think from what you've said, that sulfate has been oxidized to sulfite. Please tell me if that's right.

OpenStudy (australopithecus):

well you will be given a periodic table right?

OpenStudy (australopithecus):

electronegativity trend is pretty easy to remember

OpenStudy (jotopia34):

No, we won't be given one. or a table of half cell reactions

OpenStudy (jotopia34):

Can you just tell me if what i am thinking based off what you said is right?

OpenStudy (jotopia34):

I'm saying that sulfate is the species oxidized.

OpenStudy (australopithecus):

were you required to memorize the first few rows of elements?

OpenStudy (australopithecus):

Actually you can just guess which is which by looking at it without doing the calculations

OpenStudy (jotopia34):

also, when you say which one is more electronegative, that trips me up because sulfate and sulfite are made up of the same elements

OpenStudy (australopithecus):

http://chemwiki.ucdavis.edu/ @api/deki/files/1190/Electronegativity_Trend_IK.png?size=bestfit&width=650&height=350&revision=1

OpenStudy (australopithecus):

electronegativity is the affinity an element has for an electron

OpenStudy (jotopia34):

Oxygen I believe is more electronegative. so there is more tendency to accept where there is more oxygen. Yes?

OpenStudy (jotopia34):

or does it mean there is more to give away, analogous to having more "money" so you can donate...

OpenStudy (australopithecus):

This better exemplifies the trend, http://www.chem.umass.edu/~whelan/genchem/whelan/class_images/111_Electronegativity_Table.jpg Sulphur has less electronegativity than oxygen, that means electrons are going to hang around oxygen more than they will sulphur See my leo the lion says Grr thing to figure out which compound is oxidized and which is reduced

OpenStudy (australopithecus):

electronegativity is fundamentally important in chemistry it explains bonding

OpenStudy (australopithecus):

Sort of kind of

OpenStudy (jotopia34):

so can I assume you will not tell me if I'm right? That sulfate is oxidized species?

OpenStudy (australopithecus):

I can check one second

OpenStudy (jotopia34):

AHA! So you are not sure either. funny.

OpenStudy (australopithecus):

Oxidation state increases means loss of electrons means SO4 is oxidized relative to SO3 which is reduced because it decreases in oxidation state

OpenStudy (australopithecus):

Just look at LEO the lion says GER and knowing that SO4 has a higher oxidation state than SO3 you should be able to figure out the answer for yourself

OpenStudy (australopithecus):

Now I'm going to go back to studying because you are not a very appreciative person

OpenStudy (australopithecus):

So SO3 is oxidized to SO4, thus SO3 is the reducing agent SO4 is reduced to SO3, thus SO4 is the oxidizing agent

OpenStudy (australopithecus):

I hope you can follow what I wrote, I'm simply out of practice with these types of questions and also I have 4 exams to write this week so my brain is a little distracted. Also it is pretty offensive that when I just gave you all the information you think that I cannot answer it. I'm doing you a favour and you are antagonizing me. Anyways this is the last time I help you.

OpenStudy (jotopia34):

You could have saved me time and said you don't know for sure.

OpenStudy (australopithecus):

ugh I do know for sure I just gave you all the rules to figure it out yourself then I told you the answer.

OpenStudy (australopithecus):

The fact that you think I dont know what I'm talking about is actually pretty funny. Anyways I'm done wasting my time with you